Five hundered tickets were sold for a school play. General tickets were $4 each, while student tickets were $3 each. If the total sales were $1720, how many of each type of ticket were sold?
Answer by mananth(16949)

General tickets x numbers
Student tickets y numbers

1 x + 1 y = 500 .............1
4 x + 3 y = 1720 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 1
-3 x -3 y = -1500
4 x 3 y = 1720
Add the two equations
1 x = 220
/ 1
x = 220
plug value of x in (1)
1 x + 1 y = 500
220 + 1 y = 500
1 y = 500 -220
1 y = 280
y = 280
General tickets 220 numbers
Student tickets 280 numbers